OpenStudy (anonymous):

60 is 15% of what number?

OpenStudy (hoblos):

60 = 0.15x x=60/0.15=400

OpenStudy (anonymous):

hoblos is right .15x is 60 1% is 60/15 = 4 100% is 400

OpenStudy (anonymous):

So you divide 0.15 by 60?